A sailboat is traveling north at 10km/h relative to the water the water is flowing north at 5km/h what is the velocity of the boat releative to ground

Grade:9

2 Answers

Arun
4 years ago
Dear student
 
for angle take the dot product of above two
 
Hope it helps
 
 
 
Regards
Arun (askIITians fprum expert)
Khimraj
4 years ago
let v=the velocity of the boat relative to the ground
let x=the distance covered on the ground by the boat during time t
let vb= the velocity of the boat relative to the water
let vr=the velocity of the water.
then
v=x/t
and v= the velocity of the boat and the water, since they are both in the same direction.
So v=vb+vr
and the distance covered relative to the ground would be:
x=(vb+vr)*t